# SweetTokenizer: Semantic-Aware Spatial-Temporal Tokenizer for Compact Visual Discretization

### v2

@@ Abstract
This paper presents the Semantic-aWarE spatial-tEmporal Tokenizer (SweetTokenizer), a compact yet effective discretization approach for vision data. Our goal is to boost tokenizers’ compression ratio while maintaining reconstruction fidelity in the VQ-VAE paradigm. Firstly, to obtain compact latent representations, we decouple images or videos into spatial-temporal dimensions, translating visual information into learnable querying spatial and temporal tokens through a Cross-attention Query AutoEncoder (CQAE). Secondly, to complement visual information during compression, we quantize these tokens via a specialized codebook derived from off-the-shelf LLM embeddings to leverage the rich semantics from language modality. Finally, to enhance training stability and convergence, we also introduce a curriculum learning strategy, which proves critical for effective discrete visual representation learning. SweetTokenizer achieves comparable video reconstruction fidelity with only 25% of the tokens used in previous state-of-the-art video tokenizers, and boost video generation results by 32.9% w.r.t gFVD. When using the same token number, we significantly improves video and image reconstruction results by 57.1% w.r.t rFVD on UCF-101 and 37.2% w.r.t rFID on ImageNet-1K. Additionally, the compressed tokens are imbued with semantic information, enabling few-shot recognition capabilities powered by LLMs in downstream applications.
@@ 1 Introduction
Figure 1: Illustration of our framework. We build a compact visual latent space by reducing token count and leveraging semantic text embedding. The encoded tokens can be applied to downstream tasks, such as generation and understanding.
@@
Visual tokenizers [41, 6, 45, 43, 12, 53, 55] are emerging as essential components in the field of modern computer vision models, particularly in the generation [12, 55, 50] and understanding [44, 18, 36, 28, 47] of vision data. These tools convert visual inputs into discrete tokens, capturing essential temporal and spatial features that facilitate advanced analysis by formulating visual-related tasks as a token prediction process.
@@
Compression ratio and reconstruction fidelity are vital criteria for evaluating a tokenizer. However, recent visual tokenizers, especially video tokenizers [45, 43, 55] typically retain a low compression ratio. This is because visual tokens are usually derived from 2D patches [10] or 3D tubes [45, 12] which preserve location relationships (e.g., each token corresponds to a specific region of input [56]), leading to redundancy in both spatial and temporal dimensions. To reduce token count, we take inspiration from Q-Former [27], which compresses raw input into learnable “query tokens”. However, it is observed that directly flattening video tokens into sequence may lead to catastrophic temporal information loss. Therefore, we explore a spatial-temporal decoupled approach to perform tokenization.
@@
A higher compression ratio generally makes reconstruction more challenging. To complement visual information during compression, we follow [29, 54, 57] to introduce pretrained language embeddings as the latent codebook, leveraging the powerful semantic representation capabilities of large language models (LLM) [1, 38]. However, previous works primarily focus on image modality, overlooking the relationships between text and motion in video domain.
@@
To address existing limitations, we propose SweetTokenizer – Semantic-aWarE spatial-tEmporal Tokenizer – as illustrated in Figure 1. First, considering the heterogeneous redundancy in static images and dynamic frames, we propose the Cross-attention Query AutoEncoder (CQAE) to compress spatial and temporal information into separate learnable queries. Different from previous works [56, 27], our findings indicate that coupling the compression of spatiotemporal information increases the difficulty for the decoder to learn the motion information of the same pixel across consecutive frames. Thus, taking the decoupled spatial and temporal queries as inputs, we devise a strategy of spatial decoding followed by temporal decoding to achieve a separate reconstruction of the spatial and temporal dimensions of visual information. Additionally, the decoupled spatiotemporal reconstruction approach allows for pre-training with a large amount of image data, enhancing the model spatial representation. Second, to integrate the semantic information inherent in large language model (LLM), we design specialized codebooks tailored for spatial and temporal compression addressing the differences in semantic representation between spatial and temporal information. Specifically, we design two language-based codebooks based on the part of speech, using nouns and adjectives for spatial static information and verbs and adverbs for temporal motion information. By incorporating language-based codebooks, the learnable compressed queries can also be easily adapted to downstream visual understanding tasks by in-context learning of LLM. Third, we introduce the curriculum learning [2] mechanism by training SweetTokenizer in three stages to achieve stable convergence. We initially pre-train the spatial CQAE with image data, then joint training of the spatial and temporal CQAE with video data. The first two stages are supervised by the commitment loss of the codebook and the cross-entropy loss of the proxy code [56] from the pretrained tokenzier. In the last stage, we train the whole network and introduce the spatiotemporal decoder to reconstruct the original video.
Exhaustive experiments demonstrate the effectiveness of SweetTokenizer. Compared with SOTA methods [45], SweetTokenizer achieves comparable rFVD on the UCF-101 [34] while utilizing only 25% of the tokens. Notably, it enhances the gFVD metric from 191 to 128. At an equivalent compression rate on UCF-101, SweetTokenizer reduces rFVD from 42 to 18 and gFVD from 191 to 154. On the ImageNet-1k [8], SweetTokenizer demonstrates a substantial improvement in rFID, decreasing it from 0.59 to 0.37.
@@
  • We introduce SweetTokenizer, a cutting-edge visual tokenizer that achieves decent reconstruction fidelity with a high compression ratio via spatial-temporal decoupling and cross-attention query autoencoder, reaching a “sweet spot” between compression and fidelity.
  • We leverage a semantic-enhanced latent codebook to utilize the off-the-shell representation capability of LLM embeddings, which improves reconstruction quality and facilitates downstream video understanding tasks.
  • We design a progressive training schedule following the curriculum learning mechanism, which ensures better convergence of visual tokenizers.
  • We perform extensive experiments to verify the effectiveness of SweetTokenizer, which exhibits the state-of-the-art performance on video reconstruction, image reconstruction, and class-conditional video generation tasks, leading by a large margin of 57.1%, 37.2%, and 32.9%.
Figure 2: Main pipeline. Our semantic-aware spatial-temporal tokenizer (SweetTokenizer) is trained in a vector-quantized variational autoencoder (VQ-VAE) style. We decouple video into spatial-temporal domains, then encode/decode video via the proposed novel cross-attention query autoencoder (CQAE), which significantly reduces the token number from 5,120 to 1,280 (Section 3.2.2). We further enhance our model with a language-based latent codebook, improving reconstruction fidelity and adding extra semantics (Section 3.2.3). A curriculum learning strategy is adopted to facilitate our training in three stages: Stage 1 and Stage 2 use loss marked purple; Stage 3 uses loss marked blue (Section 3.2.4). The quantized spatial-temporal latent queries $\hat{\Omega}_{s},\hat{\Omega}_{t}$ are used for downstream tasks.
@@ 2.1 Visual Tokenizer With Vector Quantization
However, the above methods can only process image data. For video modality, TATS [12] used 3D-CNN to encode video patches and adopted sliding windows to deal with long-term relations. CViViT [43] used ViT [10] structure to encode spatial patches and then adopted a causal transformer to model temporal information. OmniTokenizer [45] and MAGVIT [53, 55] adopted similar transformer architecture and introduced image pre-training to improve video tokenizer. In this paper, we inherit the popular spatial-temporal decomposition design for video data. Moreover, we significantly increase the token compression ratio with a novel stacked cross-attention query autoencoder, which is inspired by Q-Former [27]’s token compression mechanism.
@@ 2.2 Language-based Latent Codebook
The codebooks learned by vanilla VQ-VAEs are not interpretable with lexical meanings. Therefore, many works attempt to utilize pretrained language models embedding codebooks to enhance semantics. LQAE [29] replaced the visual codebook with frozen word embeddings from BERT [9]. SPAE [54] quantized image latent space in a pyramid structure to preserve semantic information from low-level to high-level. It also used large language model (LLM) codebook [7] so that the encoded image token can be directly adapted to visual understanding tasks through in-context learning [3] ability of LLM. We follow this evaluation pipeline for few-shot classification in our paper. V2L-Tokenizer [59] utilized CLIP [31] pretrained encoder and injected a learnable projector to align visual-text latent space implicitly. VQCT [57] replaced the projector with graph convolution networks [24] to consider the relationship between vocabularies. Furthermore, De-Diffusion [48] directly encoded image into plain text as latent space interface and decodes back through a text-to-image (T2I) diffusion model [33]. However, none of these works dives deeply into the codebook design for video modality. Therefore, we propose splitting the codebook according to the video’s spatial-temporal attribute, which uses nouns & adjectives for spatial information, and verbs & adverbs for temporal information. This design helps us better align visual-motion-text semantics.
@@ 3.1 Preliminary
A typical visual vector-quantization (VQ) model [12, 45, 53, 55] contains three parts: encoder $\mathcal{E}$, decoder $\mathcal{D}$ and latent quantizer $\mathcal{Q}$. Take video modality as an example, given a video input $x\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ represents the temporal length and $H\times W$ denotes spatial resolution, encoder $\mathcal{E}(x)$ projects it into latent space $\mathbb{Z}\in\mathbb{R}^{N\times D}$, where $D$ is latent dimension and $N$ is token number. A quantizer $\mathcal{Q}$ is constructed in this latent space $\mathbb{Z}$ by querying the nearest neighbor in codebook $C\in\mathbb{R}^{L\times D}$, where $L$ is codebook size. Then $\mathcal{D}$ decodes latent space back to pixel space and applies self-supervised reconstruction loss:
@@
We aim to learn a more compact latent space $\mathbb{Z}$ while maintaining the reconstruction fidelity. To this end, we design a semantic-enhanced spatial temporal decoupled tokenizer (SweetTokenizer) to address the issue of token compression (Figure 2). The main component is Cross-attenion Query AutoEncoders (CQAE), which aggregates information from disentangled spatial-temporal dimensions into learnable queries, achieving a 4$\times$ compression ratio improvement over previous video modality tokenizers. To enhance latent space representation and reconstruction fidelity, we incorporate compact text embeddings and refine vocabularies by appearance/motion attributes for better semantic alignment. Finally, we propose a progressive training schedule to stabilize convergence.
3.2 Spatial-Temporal Tokenization
@@ 3.2.1 Patchify
Given a video frame sequence $x\in\mathbb{R}^{T\times H\times W\times 3}$, we select the first frame $x_{0}$ as a reference for spatial information, the remaining $T-1$ frames $x_{1:T}$ for temporal information, following the strategy in [45]. We apply two patch kernels with shapes $p_{h}\times p_{w}$ and $p_{t}\times p_{h}\times p_{w}$ to $x_{0}$ and $x_{1:T}$ separately, generating $v_{s}\in\mathbb{R}^{1\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ and $v_{t}\in\mathbb{R}^{\frac{T-1}{p_{t}}\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ as inputs for transformer-based autoencoder, where $v_{s}$ contains spatial information, $v_{t}$ contains temporal information, with $D$ representing the hidden embedding size. In practice, for a video with 17 frames and a resolution of $256\times 256$, we set $(p_{t},p_{h},p_{w})$ to (4, 8, 8), thus patchify frames into $v_{s}$ with shape $1\times 32\times 32$ and $v_{t}$ with shape $4\times 32\times 32$. In Figure 2, we use $t=\frac{T-1}{p_{t}}=4$ to denote $v_{t}$’s length.
Figure 3: Module structure of Cross-attention Query AutoEncoder (CQAE). Information is transported from $K,V$ head to $Q$ head through stacked CQAE transformer layers. The enlarged view illustrates the detailed architecture.
3.2.2 Cross-attention Query AutoEncoder (CQAE)
So far, the raw video has been split into patches with a fixed ratio $p_{t}\times p_{h}\times p_{w}$. However, many works [56, 27, 37] have demonstrated that visual information remains redundant spatially and temporally, even after tokenization. To select “key words” from these tokens, we adopt a strategy similar to Q-Former [27], compressing patch sequence into a fixed number of query tokens via cross-attention interactions. As an innovation, we recursively inject these cross-attention query modules into transformer-based autoencoder to transfer information, forming our CQAE module. Moreover, we show in experiments that naively flattening patches into a long 1D sequence is unsuitable for video modality since it may lose temporal motion information in consecutive frames. Therefore, we introduce two autoencoders, $CQAE_{s}$ and $CQAE_{t}$, along with two sets of latent query token embeddings $\Omega_{s}\in\mathbb{R}^{N_{s}\times D}$ and $\Omega_{t}\in\mathbb{R}^{N_{t}\times D}$, for summarizing spatial and temporal information, where $N_{s}$ and $N_{t}$ control the latent token number. We set $N_{s}$ to 256 and $N_{t}$ to 1,024, resulting in 1,280 tokens in total.
During the encoding stage, we pass $v_{s}$ into $CQAE_{s}$ and $v_{t}$ into $CQAE_{t}$. Take $CQAE_{s}$ as example, as shown in Figure 3, $v_{s}$ is first reshaped from $1\times 32\times 32$ into $1\times 1024$, then passed through $K$-layer self-attention transformer to produce $K$ intermediate hidden units $[h(v_{s}^{1}),h(v_{s}^{2}),...,h(v_{s}^{K})]$. Meanwhile, the latent query token $\Omega_{s}$ is also passed through self-attention transformers, except that each hidden unit $h(\Omega_{s}^{i})$ is served as query tensor for cross-attention layer with $h(v_{s}^{i})$. Therefore, we gradually transfer spatial information from $v_{s}$ into $\Omega_{s}$ through interactions between hidden units $h(v_{s}^{i})$ and $h(\Omega_{s})$, reducing the spatial token number from 1,024 to 256.
Similarly, $v_{t}$ is processed with $CQAE_{t}$. Instead of directly using $v_{t}$, we employ frame-wise residual $\Delta v_{t}$ to generate temporal hidden units $[h(\Delta v_{t}^{1}),h(\Delta v_{t}^{2}),...,h(\Delta v_{t}^{K})]$, as [18] found that frame motion residuals can be more effectively represented. Notably, we reshape $\Delta v_{t}$ from $4\times 32\times 32$ into $1024\times 4$, and $\Omega_{t}$ from $1024$ into $1024\times 1$ to align the transformer inputs shape. Thus, the temporal information is embedded into $\Omega_{t}$ from $v_{t}$ using the same cross-attention querying approach along the temporal axis. $\Omega_{s}$ and $\Omega_{t}$ are quantized into $\hat{\Omega}_{s}$ and $\hat{\Omega}_{t}$, which will be explained in Section 3.2.3 later.
During the decoding stage, we use mask tokens $\Omega_{m}\in\mathbb{R}^{1\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ to receive information from $\hat{\Omega}_{s}$ and $\hat{\Omega}_{t}$, similar to the mask token prediction tasks in [9, 14, 37, 56]. As shown in Figure 3, yellow blocks denote that the tensor contains spatial visual information, while grey blocks indicate that the tensor is information-free. We randomly initialize these mask tokens $\Omega_{m}$ and use them as a “canvas” to reconstruct the raw inputs by receiving visual information from $\hat{\Omega}_{s}$. The decoding stage consists of two steps. Firstly, the mask token $\Omega_{m}$ starts with shape $32\times 32$, which is reshaped into $1\times 1024$ to interact spatially with $\hat{\Omega}_{s}$. Each decoder layer has the same structure as the encoder layer, except that the query is $h(\Omega_{m}^{i})$ and the key/value is $h(\hat{\Omega}_{s}^{i})$. Afterwards, $\Omega_{m}$ is expanded from $1\times 1024$ into $5\times 1024$, in order to interact temporally with $\hat{\Omega}_{t}$. The final output $\Omega_{m}$ is a mixture of yellow and green block as illustrated in Figure 2, indicating that it simultaneously contains spatial and temporal information.
3.2.3 Language-based Latent Codebook

### v3

# SweetTok: Semantic-Aware Spatial-Temporal Tokenizer for Compact Video Discretization

@@ Abstract
This paper presents the Semantic-aWarE spatial-tEmporal Tokenizer (SweetTok), a novel video tokenizer to overcome the limitations in current video tokenization methods for compacted yet effective discretization. Unlike previous approaches that process flattened local visual patches via direct discretization or adaptive query tokenization, SweetTok proposes a decoupling framework, compressing visual inputs through distinct spatial and temporal queries via Decoupled Query AutoEncoder (DQAE). This design allows SweetTok to efficiently compress video token count while achieving superior fidelity by capturing essential information across spatial and temporal dimensions. Furthermore, we design a Motion-enhanced Language Codebook (MLC) tailored for spatial and temporal compression to address the differences in semantic representation between appearance and motion information. SweetTok significantly improves video reconstruction results by 42.8% w.r.t rFVD on UCF-101 dataset. With a better token compression strategy, it also boosts downstream video generation results by 15.1% w.r.t gFVD. Additionally, the compressed decoupled tokens are imbued with semantic information, enabling few-shot recognition capabilities powered by LLMs in downstream applications.
@@ 1 Introduction
Figure 1: Illustration of our framework. We build a compact visual latent space by reducing token count in a decoupled style and leveraging motion-enhanced semantic text embedding. The encoded tokens can be applied to downstream tasks, such as generation and understanding.
@@
Visual tokenizers [36, 5, 41, 38, 11, 48, 50, 39] are emerging as essential components in the field of modern computer vision models, particularly in the generation [11, 50, 45] and understanding [40, 14, 32, 23, 42] of vision data. These tools convert visual inputs into discrete tokens, capturing essential temporal and spatial features that facilitate advanced analysis by formulating visual-related tasks as a token prediction process.
@@
Compression ratio and reconstruction fidelity are vital criteria for evaluating a tokenizer. Recent visual tokenizers, especially video tokenizers [41, 38, 50] typically retain a low compression ratio. This is because visual tokens are usually derived from 2D patches [9] or 3D tubes [41, 11] which preserve location relationships (e.g., each token corresponds to a specific region of input [51]), leading to redundancy in both spatial and temporal dimensions. Most recent work LARP [39] quantizes the flattened video patches through adaptive holistic queries to achieve high compression ratio. However, it is observed that directly flattening video tokens into sequence may lead to difficulty in learning intertwined spatial temporal information resulting in low reconstruction performance. Therefore, a new compression method needs to be proposed, one that takes into account the spatiotemporal properties of video.
@@
Another issue, meanwhile, is that a higher compression ratio typically results in a greater loss of reconstruction details. To complement visual information during compression, one common strategy is to introduce pretrained language embeddings as the latent codebook [24, 49, 52] , leveraging their semantic representation capabilities. However, previous works primarily focus on image modality, overlooking the relationships between text and motion in video domain.
@@
To address existing limitations, we propose SweetTok – Semantic-aWarE spatial-tEmporal Tokenizer – as illustrated in Figure 1. Considering the heterogeneous redundancy in static images and dynamic frames, we propose the Decoupled Query AutoEncoder (DQAE) to compress spatial and temporal information into separate learnable queries. Different from previous works [51, 22, 39], our findings indicate that coupling the compression of spatiotemporal information increases the difficulty for the decoder to learn the motion information of the same pixel across consecutive frames. Thus, taking the decoupled spatial and temporal queries as inputs, we devise a strategy of spatial decoding followed by temporal decoding to achieve a separate reconstruction of the spatial and temporal dimensions of visual information. Additionally, the decoupled spatiotemporal reconstruction approach naturally allows for finetuning on image data, making our SweetTok flexible to image reconstruction task.
Furthermore, to integrate the semantic information inherent in pre-traiend language model, we design a Motion-enhanced Language Codebook (MLC) tailored for spatial and temporal compression addressing the differences in semantic representation between spatial and temporal information. Specifically, we design two language-based codebooks based on the part of speech, using nouns and adjectives for spatial static information and verbs and adverbs for temporal motion information. By incorporating language-based codebooks, the learnable compressed queries can also be easily adapted to downstream visual understanding tasks by in-context learning of LLM.
Exhaustive experiments demonstrate the effectiveness of SweetTok. Compared with vanilla video tokenizer without token compression (OmniTok[41]), SweetTok improves rFVD by 52.3% on UCF-101 [30] using only 25% of the tokens. Compared with vanilla query-based tokenizer (LARP [39]), SweetTok reduces rFVD from 35.15 to 20.46 and gFVD from 99 to 84, on UCF-101. By directly finetuning the decoupled spatial branch on the ImageNet-1k [7], SweetTok also demonstrates a substantial improvement in rFID, decreasing it from 0.59 to 0.37.
@@
  • We introduce SweetTok, a cutting-edge video tokenizer that achieves the state-of-the-art reconstruction fidelity with a high compression ratio via spatial-temporal decoupling and decoupled query autoencoder (DQAE), reaching a “sweet spot” between compression and fidelity.
  • We propose a motion-enhanced language codebook (MLC) to more effectively capture the action information embedded in the video modality, thereby improving reconstruction quality and supporting downstream video understanding tasks.
  • We perform extensive experiments to verify the effectiveness of SweetTok, which exhibits the state-of-the-art performance on video reconstruction, image reconstruction, and class-conditional video generation tasks, leading by a large margin of 42.8%, 37.2%, and 15.1%.
Figure 2: Pipeline overview. (a) Vanilla video tokenizers directly quantize flattened video patches. (b) Vanilla query-based tokenizers compress flattend video patches into adaptive queries. (c) SweetTok proposes decoupled query-based autoencoder (DQAE, §3.2.2). The spatial encoder quantizes the first frame’s patch embeddings, while the temporal encoder quantizes residual between consecutive frames. The spatial decoder reconstructs the first frame’s patches, replicates them $T$ times, and passes them to the temporal decoder for final information fusion and reconstruction. It also proposes motion-enhanced language codebook (MLC, §3.2.3) to complement reconstructed video information via action-related language semantics.
@@ 2.1 Visual Tokenizer With Vector Quantization
However, the above methods can only process image data. For video modality, TATS [11] used 3D-CNN to encode video patches and adopted sliding windows to deal with long-term relations. CViViT [38] used ViT [9] structure to encode spatial patches and then adopted a causal transformer to model temporal information. OmniTokenizer [41] and MAGVIT [48, 50] adopted similar transformer architecture and introduced image pre-training to improve video tokenizer. LARP [39], on the other hand, introduces to compress flattened video patches into adaptive holistic queries with the guidance of a pre-trained auto-regressive model. In this paper, we inherit the popular spatial-temporal decomposition design for video data.
@@ 2.2 Language-based Latent Codebook
The codebooks learned by vanilla VQ-VAEs are not interpretable with lexical meanings. Therefore, many works attempt to utilize pretrained language models embedding codebooks to enhance semantics. LQAE [24] replaced the visual codebook with frozen word embeddings from BERT [8]. SPAE [49] quantized image latent space in a pyramid structure to preserve semantic information from low-level to high-level. It also used large language model (LLM) codebook [6] so that the encoded image token can be directly adapted to visual understanding tasks through in-context learning [2] ability of LLM. We follow this evaluation pipeline for few-shot classification in our paper. V2L-Tokenizer [54] utilized CLIP [27] pretrained encoder and injected a learnable projector to align visual-text latent space implicitly. VQCT [52] replaced the projector with graph convolution networks [19] to consider the relationship between vocabularies. Furthermore, De-Diffusion [43] directly encoded image into plain text as latent space interface and decodes back through a text-to-image (T2I) diffusion model [29]. However, these studies primarily focus on the image modality. In this paper, we explore the design of the language codebook specifically for the video modality by splitting the codebook according to the video’s spatial-temporal attribute.
@@ 3.1 Preliminary
A typical visual vector-quantization (VQ) model [11, 41, 48, 50] contains three parts: encoder $\mathcal{E}$, decoder $\mathcal{D}$ and latent quantizer $\mathcal{Q}$ as shown in Figure 2 (a). Take video modality as an example, given a video input $x\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ represents the temporal length and $H\times W$ denotes spatial resolution, encoder $\mathcal{E}(x)$ projects it into latent space $\mathbb{Z}\in\mathbb{R}^{N\times D}$, where $D$ is latent dimension and $N$ is token number, A quantizer $\mathcal{Q}$ is constructed in this latent space $\mathbb{Z}$ by querying the nearest neighbor in codebook $C\in\mathbb{R}^{L_{c}\times D}$, where $L_{c}$ is codebook size. Then $\mathcal{D}$ decodes latent space back to pixel space and applies self-supervised reconstruction loss:
@@
Unlike traditional visual VQ models, LARP [39] quantizes flattened video patches into holistic queries $\mathbf{Q}\in\mathbb{R}^{L\times D}$ as shown in Figure 2 (b), where $L$ is the adaptive query size. As shown in Equation (2), the encoder $\mathcal{E}$ processes concatenated flattened video patches $\mathbf{E}=flatten(x)\in\mathbb{R}^{N\times D}$ and query tokens $\mathbf{Q}\in\mathbb{R}^{L\times D}$, outputting $\mathbf{Z_{Q}}\in\mathbb{R}^{L\times D}$ for quantization. The decoder $\mathcal{D}$ reconstructs the video patches $\tilde{\mathbf{x}}\in\mathbb{R}^{T\times H\times W\times 3}$ from the concatenation of learnable video queries $\mathbf{E_{Q}}\in\mathbb{R}^{N\times D}$ and the query discrete embeddings $\tilde{\mathbf{Z}}_{\mathbf{Q}}$.
$\mathbf{Z_{E}}||\mathbf{Z_{Q}}=\mathcal{E}(\mathbf{E}||\mathbf{Q}),\tilde{\mathbf{Z}}_{\mathbf{Q}}=\mathcal{Q}(\mathbf{Z_{Q}}),\tilde{x}=\mathcal{D}(\mathbf{E_{Q}}||\tilde{\mathbf{Z}}_{\mathbf{Q}}).$ (2)
However, directly compressing video from flattened patches is challenging due to intertwined redundant temporal data and low-level spatial content, leading to sub-optimal performance. Thus, we designed SweetTok to balance reconstruction performance with a high compression ratio. Details are elaborated in Section 3.2.
3.2 Decoupled Spatial-Temporal Tokenization
As noted in Section 3.1, directly quantizing video data via flattened patches hampers model learning due to intertwined redundant temporal and complex spatial information. Thus, we propose separately quantizing spatial and temporal dimensions before combining them to reconstruct the video, following the divide-and-conquer principle. The decoupling strategy enables high compression while ensuring higher fidelity. The main pipeline is shown in Figure 2 (c).
@@ 3.2.1 Patchify
Given a video frame sequence $x\in\mathbb{R}^{T\times H\times W\times 3}$, we select the first frame $x_{1}$ as a reference for spatial information, the remaining $T-1$ frames $x_{2:T}$ for temporal information, following the strategy in [41]. We apply two patch kernels $\mathcal{P}_{s},\mathcal{P}_{t}$ with shapes $p_{h}\times p_{w}$ and $p_{t}\times p_{h}\times p_{w}$ to $x_{1}$ and $x_{2:T}$ separately, generating $v_{s}\in\mathbb{R}^{1\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ and $v_{t}\in\mathbb{R}^{\frac{T-1}{p_{t}}\times\frac{H}{p_{h}}\times\frac{W}{p_{w}}\times D}$ shown below:
$v_{s}=\mathcal{P}_{s}(x_{1}),v_{t}=\mathcal{P}_{t}(x_{2:T})$ (3)
$v_{s}$ and $v_{t}$ are inputs for transformer-based autoencoder, where $v_{s}$ contains spatial information, $v_{t}$ contains temporal information. In practice, for a video with 17 frames and a resolution of $256\times 256$, we set $(p_{t},p_{h},p_{w})$ to (4, 8, 8), thus patchify frames into $v_{s}$ with shape $1\times 32\times 32$ and $v_{t}$ with shape $4\times 32\times 32$. We use $t=\frac{T-1}{p_{t}}=4$ to denote $v_{t}$’s length.
3.2.2 Decoupled Query AutoEncoder (DQAE)
To decouple spatial and temporal dimensions, we need to compress video patches separately along each dimension. Inspired by [51, 39] and Q-Former [22], we compress each dimension into an adaptive query tokens via cross-attention interactions. As an innovation, we recursively inject these cross-attention query modules into transformer-based autoencoder to transfer information, forming our DQAE module shown in the right gray box of Figure 2 (c). For the rest of our paper, we use $\mathcal{E}_{DQAE}$ and $\mathcal{D}_{DQAE}$ as the encoder and decoder of the $DQAE$ module for simplicity.
Spatial Tokenization.
We observe that for most video parts, the first frame holds the most spatial information, so we use $v_{s}$ as the input to $\mathcal{E}_{DQAE_{s}}$ for quantization shown below:
$\mathbf{Z}_{\mathbf{Q_{s}}}=\mathcal{E}_{DQAE_{s}}(\mathbf{Q_{s}},v_{s}),\tilde{\mathbf{Z}}_{\mathbf{Q_{s}}}=\mathcal{Q}_{MLC}(\mathbf{Z}_{\mathbf{Q_{s}}}),$ (4)
where $\mathbf{Q_{s}}\in\mathbb{R}^{L_{spatial}\times D}$ are the learnable spatial query embeddings and $\mathbf{Z_{\mathbf{Q_{s}}}}\in\mathbb{R}^{L_{spatial}\times D}$ are the output embeddings encoding information from the first frame patches $v_{s}$. $\tilde{\mathbf{Z}}_{\mathbf{Q_{s}}}\in\mathbb{R}^{L_{spatial}\times D}$ are the quantized spatial token embeddings and $\mathcal{Q}_{MLC}$ stands for Motion-enhanced Language Codebook (MLC) quantizer which will be elaborated in the following section. After quantization, we inject our informative spatial token embeddings $\tilde{\mathbf{Z}}_{\mathbf{Q_{s}}}$ into a learnable spatial patch queries $\mathbf{Q}_{v_{s}}$ through decoder $\mathcal{D}_{DQAE_{s}}$ as below:
$\tilde{v}_{s}=\mathcal{D}_{DQAE_{s}}(\mathbf{Q}_{v_{s}},\tilde{\mathbf{Z}}_{\mathbf{Q_{s}}}),$ (5)
where $\tilde{v}_{s}$ are the reconstructed first frame video patches. The temporal component which will be stated in the following, combined with $\tilde{v}_{s}$ is used for the final video reconstruction. We set $L_{spatial}=256$ in real implementation.
Temporal Tokenization.
It is observed that for video data, there is much redundancy along temporal dimension. It motivates us to employ frame-wise residual $\Delta v_{t}=(\Delta v_{t}^{1},\Delta v_{t}^{2},...,\Delta v_{t}^{k})_{k=\frac{H}{p_{h}}\times\frac{W}{p_{w}}}$, where $\Delta v_{t}^{i}=v_{s}^{i}-v_{t}^{i}$, for tokenization. We use the first frame of the video for frame-wise residual because the spatial tokenization phase reconsturct it. Then, the residual $\Delta v=(\Delta v_{1},\Delta v_{2},...,\Delta v_{t})_{t=\frac{T-1}{p_{t}}}$ is input to $\mathcal{E}_{DQAE_{t}}$ for temporal compression, as shown below:
$\mathbf{Z}_{\mathbf{Q_{t}}}=\mathcal{E}_{DQAE_{t}}(\mathbf{Q_{t}},\Delta v),\tilde{\mathbf{Z}}_{\mathbf{Q_{t}}}=\mathcal{Q}_{MLC}(\mathbf{Z}_{\mathbf{Q_{t}}}),$ (6)
where $\mathbf{Q_{t}}\in\mathbb{R}^{L_{temporal}\times D}$ are the learnable temporal query embeddings and $\mathbf{Z_{\mathbf{Q_{t}}}}\in\mathbb{R}^{L_{temporal}\times D}$ are the output embeddings encoding information of the frame-wise residual. $\tilde{\mathbf{Z}}_{\mathbf{Q_{t}}}\in\mathbb{R}^{L_{temporal}\times D}$ are the quantized frame-wise temporal residual token embeddings. In practical implementation, $L_{temporal}=1024$. After reconstructing the first frame patches, we recover the entire video patches by combining $\tilde{v}_{s}$ with the frame-wise quantized residual $\tilde{\mathbf{Z}}_{\mathbf{Q_{t}}}$. We tiled $\tilde{v}_{s}$ for $t$ times and sent it to $\mathcal{D}_{DQAE_{t}}$ shown below:
$\tilde{v}=\mathcal{D}_{DQAE_{t}}([\tilde{v}_{s}||\cdots||\tilde{v}_{s}],\tilde{\mathbf{Z}}_{\mathbf{Q_{t}}}),$ (7)
where $\tilde{v}$ is the reconstructed video patches. The vectors $\tilde{v}$ reside in the latent space, necessitating the use of a “pixel decoder” $\mathcal{D}_{pixel}$ to reconstruct the video data, as illustrated below:
$\tilde{x}=\mathcal{D}_{pixel}(\tilde{v}).$ (8)
The whole $DQAE$ is supervised by the reconstruction loss $\mathcal{L}_{rec}$ containing a $L_{2}$ loss, a LPIPS perception loss: $\mathcal{L}_{Lpips}$, a quantizer loss: $\mathcal{L}_{vq}$ and a GAN loss: $\mathcal{L}_{g}$ following the principle [10].
3.2.3 Motion-enhanced Language Codebook (MLC)